Grade 12Engineering Entrance Exams

an object initially at rest, starts moving with a uniform accleration . its speed at the end of 2nd second is 4m/s and at the end of 7th second is 14m/s. what is the distance travelled during the interval of observation

nikitha
10 Years agoGrade 12

1 Answer

Riddhish Bhalodia
10 Years ago
Let the uniform acceleration be a
so till t=2 we have
v = u + at
4 = a x 2
hence a = 2
Now we need to find the distance between 2ndsecond and the 7thsecond
so we use
v^2 - u^2 = 2as
so we have
196 – 16 = 2 x 2 x s
s = 45m

question mark
Hello,
 
Can we learn and prepare for MHT-CET online ?
engineering entrance exams

Last Activity: 4 Years ago